OpenStudy (anonymous):

What is the value of the expression x + (–y) when x = 1.2 and y = –10.7? A. 9.5 B. 11.9 C. –11.9 D. –9.5 I think that it is D

OpenStudy (anonymous):

1.2+10.7=

OpenStudy (anonymous):

B

OpenStudy (anonymous):

it is -10.7

OpenStudy (anonymous):

x + (-(-10.7)) = 11.9

OpenStudy (anonymous):

it's because you have x + (-y) when put plug -10.7 in to y you have -1*-10.7 which makes it +10.7

OpenStudy (anonymous):

in other words it just cancels out the negative.

jimthompson5910 (jim_thompson5910):

\[\Large \color{blue}{x = 1.2}\] \[\Large \color{red}{y = -10.7}\] ------------------------------------------------------- \[\Large \color{blue}{x} + (-\color{red}{y})\] \[\Large \color{blue}{x} - (\color{red}{y})\] \[\Large \color{blue}{1.2} - (\color{red}{-10.7})\] \[\Large 1.2 + 10.7\] \[\Large 11.9\] Notice how the two negatives canceled to form a positive

OpenStudy (anonymous):

no that is just purenlesys this is what makes is a positive 10 | 10.7|

OpenStudy (anonymous):

sorry i meant | -10.7| = 10.7 or |10.7| = -10.7

OpenStudy (anonymous):

no that is just purenlesys this is what makes is a positive 10 | 10.7|

OpenStudy (anonymous):

|number| is always \[\ge 0\]

OpenStudy (anonymous):

sorry i meant | -10.7| = 10.7 or |10.7| = -10.7

OpenStudy (anonymous):

you seem to be repeating yourself.

OpenStudy (anonymous):

Any one But lily2913 have any thing to help P.S. @lily2913 thank you for your help

jimthompson5910 (jim_thompson5910):

So the problem is actually \(\Large x+|y|\) ?

OpenStudy (anonymous):

yes x = 1.2 and y = –10.7?

jimthompson5910 (jim_thompson5910):

you are correct in saying that `|-10.7| = 10.7`

jimthompson5910 (jim_thompson5910):

so, `x+|y| = 1.2 + |-10.7| = 1.2 + 10.7`

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

Hello

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you so MUCH

jimthompson5910 (jim_thompson5910):

you're welcome